OpenStudy (anonymous):

Use the properties of exponents to explain why \(64^{\frac{1}{2}}\) is called the \(\sqrt{64}\).

OpenStudy (anonymous):

@Data_LG2

OpenStudy (anonymous):

i know how it is, but i am stuck on how to explain

OpenStudy (anonymous):

\(x^{\frac{1}{2}}\) = \(\sqrt{x}\)

OpenStudy (anonymous):

see I have this so far Looking at the rational exponent property, if you have a number (x) with an exponent of a fraction (1/2 for example), you are able to change it to a radical, and with the denominator being 2 and the numerator being 1, x^(1/2) = sqrt(x), and the same goes for 64^(1/2). i feel like i need to explain it better

OpenStudy (anonymous):

@SolomonZelman @texaschic101 @hartnn

OpenStudy (anonymous):

@Luigi0210 @pooja195 @Nnesha

Nnesha (nnesha):

example can be \[\huge\rm \sqrt[n]{x^m} = x^\frac{ m}{ n }\]

OpenStudy (anonymous):

yeah i know

OpenStudy (anonymous):

i just have no clue how to explain it right

Nnesha (nnesha):